the area of a rectangular field is 300M. its perimeter is 70M.find the width and length of this rectangle
Answer by stanbon(75887)

the area of a rectangular field is 300M. its perimeter is 70M.find the width and length of this rectangle
---
Equations:
LW = 300M
L+W = 35
----
Substitute for "L" and solve for "W":
(35-W)W = 300
-W^2+35W-300
---
W^2-35W+300 = 0
---
Factor:
(W-20)(W-15)=0
---
If W = 20, L = 15
If W = 15, L = 20
==================================
Cheers,
Stan H.